OpenStudy (anonymous):

Find the integrating factor and use it to find the general solution for the following differential equation: ty'+y = -5tcos(8t)

OpenStudy (anonymous):

@Loser66 @wio i would have to divide out the t correct

OpenStudy (loser66):

@wio

OpenStudy (loser66):

to me, Yes

OpenStudy (anonymous):

so its y'+y = -5cos(8t)

OpenStudy (loser66):

the middle term is \(\dfrac{1}{t} y\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Look again at the left hand side \[\Large ty'+y=\frac{d(yt)}{dt} \]

OpenStudy (anonymous):

So you're already set :)

OpenStudy (anonymous):

wait what...

OpenStudy (anonymous):

Finding an integration factor isn't necessary, but if you do, it will turn out to be \(t\) which you need to multiply your DE through in normal standard form, so it will turn out to be the same equation as you have already written up there.

OpenStudy (anonymous):

@RichG1, the left side is already the derivative of another expression. All that's left is to integrate.

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

Dividing out by \(t\) yields \[y'+\frac{1}{t}y=-5\cos8t\] The integrating factor would be \(\mu(t)=e^{\int1/t~dt}=e^{\ln|t|}=t\). Multiplying by \(\mu\) yields the same original equation: \[ty'+y=-5t\cos8t\] What we're all trying to say is that the integrating factor-finding is not needed here, because you can proceed with solving for \(y\) right away.

OpenStudy (anonymous):

so what im getting here is \[e ^{1/t} \int\limits_{} -5\cos8t e ^{1/t}dt +ce ^{1/t}\]

OpenStudy (anonymous):

No, the integrating factor was \(t\), not \(e^{1/t}\). Given \[ty'+y=-5t\cos8t\] You have that the left side is derivative of a product: \[\frac{d}{dt}[ty]=-5t\cos8t\\ ty=-5\int t\cos8t~dt\]

OpenStudy (anonymous):

ok then intergrate and devide out t

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

so -5/8 sin(8 t) + c ?

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

No, the integral wasn't done correctly. First, you have to integrate by parts: \[\begin{matrix}u=t&&&dv=\cos8t~dt\\du=dt&&&v=\frac{1}{8}\sin8t\end{matrix}\] \[\begin{align*}-5\int t\cos8t~dt&=-5\left[\frac{1}{8}t\sin8t-\frac{1}{8}\int\sin8t~dt\right]\\ &=-\frac{5}{8}t\sin8t-\frac{5}{16}\cos8t+C \end{align*}\] So you have \[ty=-\frac{5}{8}t\sin8t-\frac{5}{16}\cos8t+C\]

OpenStudy (anonymous):

Oh i tried to factor out the t

OpenStudy (anonymous):

i dont think this is correct when i factor out the t i get \[\frac{-5 \cos8 t+2 t \sin8t }{ 16 t } \]